# ELITR-Bench: A Meeting Assistant Benchmark for Long-Context Language Models

### v2

@@ Abstract
Research on Large Language Models (LLMs) has recently witnessed an increasing interest in extending models’ context size to better capture dependencies within long documents. While benchmarks have been proposed to assess long-range abilities, existing efforts primarily considered generic tasks that are not necessarily aligned with real-world applications. In contrast, our work proposes a new benchmark for long-context LLMs focused on a practical meeting assistant scenario. In this scenario, the long contexts consist of transcripts obtained by automatic speech recognition, presenting unique challenges for LLMs due to the inherent noisiness and oral nature of such data. Our benchmark, named ELITR-Bench, augments the existing ELITR corpus’ transcripts with 271 manually crafted questions and their ground-truth answers. Our experiments with recent long-context LLMs on ELITR-Bench highlight a gap between open-source and proprietary models, especially when questions are asked sequentially within a conversation. We also provide a thorough analysis of our GPT-4-based evaluation method, encompassing insights from a crowdsourcing study. Our findings suggest that while GPT-4’s evaluation scores are correlated with human judges’, its ability to differentiate among more than three score levels may be limited.
@@ 1 Introduction
The context window of Large Language Models (LLMs) has recently undergone a significant expansion, scaling from a few thousand tokens to tens or even hundreds of thousands [8, 45, 28, 9, 4]. As a consequence, benchmarks have emerged to assess LLMs’ long-range abilities, tackling the specific challenges of Question Answering (QA) on long documents [1, 25, 3, 26, 31, 48]. However, while previous datasets focusing on long-context models offer longitudinal evaluations across different tasks, they often provide only superficial analyses of each task. The covered tasks are also often generic – e.g., questions on Wikipedia [26] – and thus not particularly suitable for realistic, focused applications. [^1] Aside from this very recent work [27] (WildBench), which offers an automated evaluation framework for assessing LLMs with complex real-world user queries.
In contrast, our work advocates for a situated evaluation of long-context LLM performance within specific, real-world scenarios. As a practical illustration, consider a meeting assistant that allows users to inquire about meetings they did not attend. Given that hour-long meeting transcripts must fit within the agent’s context window, proficient handling of long contexts is a prerequisite. This paper then introduces the first benchmark – to the best of our knowledge – for evaluating long-context LLMs on a realistic meeting assistant task. Our benchmark, named ELITR-Bench, [^2] We release the data associated to our benchmark, as well as the generations produced by compared models and their evaluation, at https://github.com/utter-project/ELITR-Bench. is built upon the meeting transcripts of the past ELITR project [33]. These transcripts have been obtained by Automatic Speech Recognition (ASR) with minimal human corrections – yielding long, noisy documents of oral nature that present unique challenges for LLMs. Our extensive experiments on ELITR-Bench with 9 recent long-context LLMs showed a gap between proprietary and open-source models that is emphasized when questions are asked sequentially within a conversation rather in a QA mode. We also provide an in-depth validation of our GPT-4-based evaluation through a crowdsourcing study, showing a high correlation with human judges’ scores. However, we also point out GPT-4’s limited ability to distinguish beyond three score levels.
@@ 3 ELITR-Bench
We build our benchmark on top of the ELITR Minuting Corpus [33]. [^5] Accessible at: https://lindat.mff.cuni.cz/repository/xmlui/handle/11234/1-4692 This corpus contains transcripts of meetings conducted in both Czech and English, along with manually crafted summaries referred to as ‘minutes’. The meeting durations range from 10 minutes to over 2 hours, with the majority lasting around one hour. Although transcripts have been corrected from ASR outputs, they still contain noise and reflect various oral language phenomena such as interjections. Each transcript is de-identified [^6] The authors ensured the removal or masking of any personally identifiable information (PII), such as names, addresses, or other details from the transcripts. Moreover, they de-identified any project or organization-related information, as its inclusion could indirectly reveal the individuals involved. and accompanied by one or multiple corresponding minutes files. However, in the benchmark described here, we only use the verbatim transcripts and exclude the minutes. In the current version of ELITR-Bench, our focus is on English meetings. Specifically, we utilized the official dev and test2 sets, consisting of 10 and 8 meetings respectively, both sourced from ELITR-English. These meetings focus on discussions related to the computer science domain, with a particular emphasis on Natural Language Processing (NLP) topics. For every meeting within this corpus, we meticulously formulated a series of questions that can be directly addressed using the corresponding transcript, and provided their corresponding ground-truth answers. We present in Appendix D (Table 13) a snippet of an ELITR meeting transcript, and showcase examples of questions and answers introduced in ELITR-Bench.
@@
The questions we defined span various types, including: Who questions, What questions (that also cover Why questions), When questions, and How many questions. Additionally, we annotated the position of the answer within the meeting transcript, categorizing it as either in the Beginning (first third), Middle (second third), End (final third), or spanning Several passages throughout the transcript. This annotation was conducted to verify the findings of Liu et al. [29], which suggest that LLMs may face challenges in processing information located in the middle of long contexts, potentially leading to performance degradation.
@@
Table 1 provides a summary of the statistics for our benchmark. In the upcoming sections, we will showcase the performance of long-context LLMs on ELITR-Bench, particularly in their ability to handle hour-long meetings – which requires processing extended contexts of more than 10K tokens on average.
@@
average [min; max]
Dev 10 141 What 59 Begin 45 11,339 [5,152; 17,410]
@@
Test 8 130 What 57 Begin 43 12,562 [4,779; 17,615]
@@ 4 Experimental setup
The evaluation on ELITR-Bench is conducted as follows. For each meeting, a prompt containing the transcript and detailing the assistant’s task is formed. Then, questions are appended to the initial prompt to drive the conversation about the corresponding meeting. We consider two ways to do this: (i) the single-turn mode, where only a single question is tackled in the conversation (i.e., the prompt is re-initialized for each new question), or (ii) the multi-turn mode, where all the questions related to a meeting are asked successively within a single conversation. Given the stand-alone nature of questions in ELITR-Bench-QA, one can adopt either the single-turn or multi-turn modes for this setting, whereas for ELITR-Bench-Conv it only makes sense to use the multi-turn mode as some questions are inter-dependent. In our evaluation methodology, given a question integrated in the aforementioned prompt, the response generated by an LLM is evaluated automatically using a GPT-4 judge, [^7] Our GPT-4 judge is based on the gpt-4-0613 checkpoint, for its cheaper cost compared to gpt-4-turbo models. Pilot experiments with different GPT-4 judges led to similar evaluation scores. following the standard practice in LLM evaluation (as discussed in Section 2). Specifically, we adopted a score rubric-based evaluation methodology [21] in which a generated response is evaluated on its proximity to the ground-truth answer, given the associated question and a score rubric that details the quality criteria expected at each score level (ranging from the lowest score of 1 to the perfect score of 10). The prompt used for the evaluation as well as our manually defined score rubric are provided in Appendix E (Figs. 5 and 6, respectively). Although our core experiment results rely on automatic LLM-based evaluation (Section 5), we further confirm the validity of this methodology against human judgement in Section 6.
@@
In our experiments on ELITR-Bench, we compared responses generated by 9 recent LLMs with long-context capabilities. We included both commercial models and open-source long-context models based on LLaMA-2 in our benchmarking:
  • GPT-3.5 and GPT-4 [34] are powerful commercial LLMs from OpenAI that have obtained state-of-the-art performance on a wide range of LLM benchmarks. We used the gpt-3.5-turbo-16k-0613 and gpt-4-1106-preview checkpoints, [^8] https://platform.openai.com/docs/models/ which respectively enable a context length of 16K and 128K tokens.
We provide more details on the compared models in Appendix A.1. Additionally, we described the search conducted to select the best configuration (including inference hyperparameters and prompt formatting) for each model in Appendix A.2.
@@ 5 Experimental results
This section describes the results of the experiments conducted on ELITR-Bench. In Section 5.1, we summarize the benchmarking results of the compared models on ELITR-Bench-QA and ELITR-Bench-Conv. Then, in Section 5.2, we analyze the impact of the question types and answer positions on the models’ performance. Finally, Section 5.3 discusses the results for questions that differ between the QA and Conv settings.
@@ 5.1 Main results
The main results of the benchmarking on ELITR-Bench are reported in Table 2. The compared models are evaluated in three settings that combine the ELITR-Bench-QA or ELITR-Bench-Conv question set with the single-turn mode (i.e., one question asked per conversation) or multi-turn mode (i.e., all questions related to one meeting asked in a single conversation). [^9] Single-turn ELITR-Bench-Conv is omitted as some questions in the Conv setting are context-dependent (i.e., rely on previous questions or answers) and thus could not be asked independently. For each of the three considered settings, we report the results on the dev set, the results on the test set, and their mean. Given the extensive cost of GPT4-based evaluation (detailed in Section 4), we performed a single seeded run for the dev set and three seeded runs for the test set. For the latter, we report the average score over the three runs. In Appendix B.1, we provide more details about the seeded runs as well as their standard deviations.
@@
Looking at the three settings in Table 2, we observe that GPT-4 clearly dominates over all other approaches with an average score that is always above 8.101010While one might argue that GPT-4 is unfairly advantaged due to the use of a GPT-4 judge, we show in Section 6.2 that the dominance of this model is observed for other evaluators as well. GPT-3.5 obtained a slightly lower average score  – around 7 – that still outperforms open-source LLMs. Among these, differences are smaller with scores close to 6 on the single-turn setting, and ranging from 4 to 6 on the multi-turn settings. Nonetheless, we can note that Vicuna-13B-v1.5 is the open-source approach that performed the most favorably overall on the three settings. Interestingly, the results in the single-turn and multi-turn modes show large discrepancies for open-source models – even when the question set is exactly the same, for ELITR-Bench-QA. This seems to indicate that open-source long-context LLMs get distracted by the previous questions and answers, which affects their performance. In contrast, GPT-4 is instead able to increase its performance between the single-turn mode and the multi-turn mode. Comparing the results of the QA and Conv settings in the multi-turn mode, we found only minimal differences. This can be explained by the small number of questions that differ between QA and Conv (16 for the dev set and 17 for the test set). In Section 5.3, we analyze the results on this subset of differentiating questions to better understand the impact of the benchmark setting (QA vs Conv) and inference mode (single-turn vs multi-turn).
@@
Table 2: Results on different ELITR-Bench settings. The reported numbers correspond to the average scores from 1 to 10 (higher is better) obtained by a GPT-4 evaluator, on a single seeded run for the dev set and 3 seeded runs for the test set. Boldface numbers correspond to the best performance among proprietary or open-source models. The results for GPT-3.5 are omitted in the multi-turn setting as the context length exceeded the 16K limit of this model.
@@ 5.2 Impact of question type and answer position
As each question in ELITR-Bench is characterized by its type (Who, What, When, How many) and answer location (Beginning, Middle, End, Several), we sought to identify whether these impact the models’ ability to answer correctly. We show in Table 3 the results restricted to each question type and answer location, obtained on the test set of ELITR-Bench-QA in single-turn mode. Due to space limitations, we aggregate the results by family of models (GPT models or LLaMA-2 models) to look for general trends among comparable models. The detailed, per-model results are available in Appendix B.2 (Table 9).
Looking at the question type results, we find that for both model families the Who questions are the easiest to answer. In contrast, the What questions were the most challenging for GPT models and the second most challenging for LLaMA-2 models. This is not surprising as What questions sometimes require complex answers that go beyond simply listing entities, dates or numbers. Interestingly, LLaMA-2 models struggled the most with the How many questions. Although the amount of such questions is very limited (8 in the test set) which calls for caution on tentative interpretations, this seems to suggest that LLaMA-2 models are notably less proficient at dealing with quantities and numbers than GPT models.
In contrast, the results by answer location in Table 3 do not seem to show any general trend. In particular, we do not notice at first glance any “lost in the middle” effect [29] which posits that information located in the middle section of long contexts is harder to access for LLMs. To further verify this, we conducted a one-tailed Welch’s t-test [43] for each model to investigate the hypothesis stating that the model’s average score for questions with middle-position answers is lower than that of other questions. We found that this hypothesis is only verified for two models: LongChat-7B-v1.5 (p-value = 0.032) and Vicuna-7B-v1.5 (p-value = 0.046) – the full results are available in Appendix B.2 (Table 10). This suggests that all models may not be affected in the same way by the location of information in the context.
Table 3: Results by question type and answer location for the GPT family (2 models) and the LLaMA-2 family (7 models) on the test set of ELITR-Bench-QA in single-turn mode. The number N below a subset indicates the corresponding subset size.
Model family Question type Answer location
@@
GPT 8.24 7.62 7.98 8.04 7.85 7.87 8.04 7.97
LLaMA-2 6.50 5.88 6.00 5.29 6.31 5.84 6.00 6.07
5.3 Results on QA/Conv differentiating questions
Similarly with what we observed in Table 2, we notice again a clear difference between GPT-4 and open-source models: the performance of the former improves (slightly) from single-turn to multi-turn, whereas the performance of the latter notably degrades. In contrast with our previous findings that showed little to no difference between the results on ELITR-Bench-QA and ELITR-Bench-Conv for the multi-turn mode, we observe this time that the average score decreases from QA to Conv for open-source models. While the difference is small, this trend is present for all open-source models except LongChat-7B-v1.5. This trend was expected as the Conv questions in this subset are more challenging to answer. However, interestingly, GPT-4 results did not show the same trend. We hypothesize that the opposite trends identified for GPT-4 and open-source models might be explained by a ‘snowballing’ effect that causes an error propagation in lower-performing open-source models and instead provides additional helpful context for GPT-4.

### v3

@@ Abstract
Research on Large Language Models (LLMs) has recently witnessed an increasing interest in extending the models’ context size to better capture dependencies within long documents. While benchmarks have been proposed to assess long-range abilities, existing efforts primarily considered generic tasks that are not necessarily aligned with real-world applications. In contrast, we propose a new benchmark for long-context LLMs focused on a practical meeting assistant scenario in which the long contexts consist of transcripts obtained by automatic speech recognition, presenting unique challenges for LLMs due to the inherent noisiness and oral nature of such data. Our benchmark, ELITR-Bench, augments the existing ELITR corpus by adding 271 manually crafted questions with their ground-truth answers, as well as noisy versions of meeting transcripts altered to target different Word Error Rate levels. Our experiments with 12 long-context LLMs on ELITR-Bench confirm the progress made across successive generations of both proprietary and open models, and point out their discrepancies in terms of robustness to transcript noise. We also provide a thorough analysis of our GPT-4-based evaluation, including insights from a crowdsourcing study. Our findings indicate that while GPT-4’s scores align with human judges, its ability to distinguish beyond three score levels may be limited.
@@ 1 Introduction
The context window of Large Language Models (LLMs) has recently undergone a significant expansion, scaling from a few thousand tokens to tens or even hundreds of thousands [9, 48, 30, 10, 5]. As a consequence, benchmarks have emerged to assess LLMs’ long-range abilities, tackling the specific challenges of Question Answering (QA) on long documents [2, 27, 4, 28, 33, 51]. However, while previous datasets focusing on long-context models offer longitudinal evaluations across different tasks, they often provide only superficial analyses of each task. The covered tasks are also often generic – e.g., questions on Wikipedia [28] – and thus not particularly suitable for realistic, focused applications. [^1] WildBench [29] is a very recent exception as it proposes an automated evaluation framework for assessing LLMs with complex real-world user queries.
In contrast, our work advocates for a situated evaluation of long-context LLM performance within specific, real-world scenarios. As a practical illustration, consider a meeting assistant that allows users to inquire about meetings they did not attend. Given that hour-long meeting transcripts must fit within the agent’s context window, proficient handling of long contexts is a prerequisite. This paper then introduces the first benchmark – to the best of our knowledge – for evaluating long-context LLMs on a realistic meeting assistant task. Our benchmark, named ELITR-Bench, [^2] We release the data associated to our benchmark, as well as the generations produced by compared models and their evaluation, at https://github.com/utter-project/ELITR-Bench. is built upon the meeting transcripts of the past ELITR project [35]. These transcripts have been obtained by Automatic Speech Recognition (ASR) with minimal human corrections – yielding long, noisy documents of oral nature that present unique challenges for LLMs. ELITR-Bench enables tracking the progress of successive model generations (e.g., GPT-3.5 vs GPT-4/4o, or LLaMA-2 vs LLaMA-3.1), as demonstrated by our extensive experiments on 12 long-context proprietary and open LLMs. Additionally, we analyzed how these models handle varying noise levels in meeting transcripts. Our GPT-4-based evaluation was validated through a crowdsourcing study, which revealed a high correlation with human judges’ scores but also a limited ability to distinguish between more than three quality levels despite using a score range of 10 points.
@@ 3 ELITR-Bench
We build our benchmark on top of the ELITR Minuting Corpus [35]. [^5] Accessible at: https://lindat.mff.cuni.cz/repository/xmlui/handle/11234/1-4692 This corpus contains transcripts of meetings conducted in both Czech and English, along with manually crafted summaries referred to as ‘minutes’. The meeting durations range from 10 minutes to over 2 hours, with the majority lasting around one hour. Although transcripts have been corrected from ASR outputs, they still contain noise and reflect various oral language phenomena such as interjections. Each transcript is de-identified [^6] The authors ensured the removal or masking of any personally identifiable information (PII), such as names, addresses, or other details from the transcripts. Moreover, they de-identified any project or organization-related information, as its inclusion could indirectly reveal the individuals involved. and accompanied by one or multiple corresponding minutes files. However, in the benchmark described here, we only use the verbatim transcripts and exclude the minutes. In the current version of ELITR-Bench, our focus is on English meetings. Specifically, we utilized the official dev and test2 sets, consisting of 10 and 8 meetings respectively, both sourced from ELITR-English. These meetings focus on discussions related to the computer science domain, with a particular emphasis on Natural Language Processing (NLP) topics. For every meeting within this corpus, we manually created a series of questions that can be directly addressed using the corresponding transcript, and provided their corresponding ground-truth answers. We present in Appendix E (Table 12) a snippet of an ELITR meeting transcript, and showcase examples of questions and answers introduced in ELITR-Bench.
@@
The questions we defined span various types, including: Who questions, What questions (which also include Why questions), When questions, and How many questions. Additionally, we annotated the position of the answer within the meeting transcript, categorizing it as either in the Beginning (first third), Middle (second third), End (final third), or spanning Several passages throughout the transcript. Table 1 provides a summary of the statistics for our benchmark.
@@
average [min, max]
Dev 10 141 What 59 Begin 45 11.3k [5.2k, 17.4k]
@@
Test 8 130 What 57 Begin 43 12.6k [4.8k, 17.6k]
@@
Noisy versions of the meeting transcripts.
To assess the robustness of long-context LLMs to noisy text, we created multiple noisy versions of the ELITR meeting transcripts by simulating different levels of ASR noise. Using a large corpus of over 500k ASR transcripts aligned with references, [^7] https://huggingface.co/datasets/google/red_ace_asr_error_detection_and_correction we generated 86,148 substitution rules where each rule consists of a token and a probability distribution over similar tokens (or an empty character) that can plausibly replace it. This extensive rule set enables us to simulate noisy transcripts with varying target Word Error Rate levels (20%, 40%, 60%, 80% and 100%). [^8] In practice, the effective Word Error Rate obtained after applying the noise injection procedure to the transcripts is generally lower than the target one – it is therefore still possible to infer some answers with a target level of 100%. We provide more details in Appendix A.
@@ 4 Experimental setup
The evaluation on ELITR-Bench is conducted as follows. For each meeting, a prompt containing the transcript and detailing the assistant’s task is formed. Then, questions are appended to the initial prompt to drive the conversation about the corresponding meeting. We consider two ways to do this: (i) the single-turn mode, where only a single question is tackled in the conversation (i.e., the prompt is re-initialized for each new question), or (ii) the multi-turn mode, where all the questions related to a meeting are asked successively within a single conversation. Given the stand-alone nature of questions in ELITR-Bench-QA, one can adopt either the single-turn or multi-turn modes for this setting, whereas for ELITR-Bench-Conv it only makes sense to use the multi-turn mode as some questions are inter-dependent. In our evaluation methodology, given a question integrated in the aforementioned prompt, the response generated by an LLM is evaluated automatically using a GPT-4 judge, [^9] Our GPT-4 judge is based on the gpt-4-0613 checkpoint, for its cheaper cost compared to gpt-4-turbo models. Pilot experiments with different GPT-4 judges led to similar evaluation scores. following the standard practice in LLM evaluation (as discussed in Section 2). Specifically, we adopted a score rubric-based evaluation methodology [23] in which a generated response is evaluated on its proximity to the ground-truth answer, given the associated question and a score rubric that details the quality criteria expected at each score level (ranging from the lowest score of 1 to the perfect score of 10). The prompt used for the evaluation as well as our manually defined score rubric are provided in Appendix F (Figs. 6 and 7, respectively). Although our core experiment results rely on automatic LLM-based evaluation (Section 5), we further confirm the validity of this methodology against human judgement in Section 6.
@@
In our experiments on ELITR-Bench, we compared responses generated by 12 LLMs with long-context capabilities (at least 32k tokens). We included both commercial models and open long-context models based on LLaMA-{2, 3.1} and Phi-3 in our benchmarking:
  • GPT-3.5, GPT-4 [36], GPT-4o;
  • LLaMA-2 models extended for long-context scenarios: LongAlpaca-{7B, 13B} [10], LongChat-7B-v1.5 [27], Vicuna-{7B, 13B}-v1.5 [12], LongAlign-{7B, 13B} [5];
  • LLaMA-3.1-8B [17];
  • Phi-3-small [1].
We provide more details on the compared models in Appendix B.1. Additionally, we describe the search conducted to select the best model configuration (including inference hyperparameters and prompt formatting) in Appendix B.2.
@@ 5 Experimental results
This section describes the results of the experiments conducted on ELITR-Bench. In Section 5.1, we summarize the benchmarking results of the compared models on ELITR-Bench-QA and ELITR-Bench-Conv. Then, in Section 5.2, we analyze the impact of the question types and answer positions on the models’ performance. Finally, Section 5.3 discusses how the noise level of the meeting transcripts influences the performance of the models.
@@ 5.1 Main results
The main results of the benchmarking on ELITR-Bench are reported in Table 2. The compared models are evaluated in three settings that combine the ELITR-Bench-QA or ELITR-Bench-Conv question set with the single-turn mode (i.e., one question asked per conversation) or multi-turn mode (i.e., all questions related to one meeting asked in a single conversation). [^10] Single-turn ELITR-Bench-Conv is omitted as some questions in the Conv setting are context-dependent (i.e., rely on previous questions or answers) and thus could not be asked independently. For each of the three considered settings, we report the results on the dev set, the results on the test set, and their mean. Given the extensive cost of GPT4-based evaluation (detailed in Section 4), we performed a single seeded run for the dev set and three seeded runs for the test set. For the latter, we report the average score over the three runs. In Appendix C.1, we provide more details about the seeded runs as well as their standard deviations.
@@
Looking at the three settings in Table 2, we observe that GPT-4 and GPT-4o dominate over all other approaches with an average score that is always above 8.111111While one might argue that GPT-4 is advantaged due to the use of a GPT-4 judge, we show in Section 6.2 that this model’s dominance is observed for other evaluators as well. GPT-3.5 obtained a slightly lower average score – around 7 – and was also beaten by the two more recent open LLMs, LLaMA-3.1-8B and Phi-3-small with LLaMA-3.1-8B coming out on top. These three models notably outperformed the LLaMA-2-based LLMs. Among the latter, differences are smaller with scores close to 6 on the single-turn setting, and ranging from 4 to 6 on the multi-turn settings. Nonetheless, we can note that Vicuna-13B-v1.5 is the LLaMA-2-based approach that performed the most favorably overall on the three settings. Interestingly, the results in the single-turn and multi-turn modes show large discrepancies for LLaMA-2-based models – even when the question set is exactly the same, for ELITR-Bench-QA. This seems to indicate that these LLMs get distracted by the previous questions and answers, which affects their performance. In contrast, GPT-4/4o is able to maintain its performance between the single-turn mode and the multi-turn mode. The same can be observed for LLaMA-3.1-8B and Phi-3-small, suggesting that recent open long-context LLMs are able to successfully handle multi-turn conversations, unlike their predecessors. Comparing the results of the QA and Conv settings in the multi-turn mode, we found only minimal differences. This can be explained by the small number of questions that differ between QA and Conv (16 for the dev set and 17 for the test set). In Appendix C.2 (Fig. 3), we further analyze the impact of the benchmark setting (QA vs Conv) and inference mode (single-turn vs multi-turn) by detailing the results restricted to this subset of differentiating questions.
@@
Table 2: Results on different ELITR-Bench settings. The reported numbers correspond to the average scores from 1 to 10 (higher is better) obtained by a GPT-4 evaluator, on a single seeded run for the dev set and 3 seeded runs for the test set. Boldface numbers correspond to the best performance among proprietary or open models. The results for GPT-3.5 are omitted in the multi-turn setting as the context length exceeded the 16k limit of this model.
@@ 5.2 Impact of question type and answer position
Table 3: Results by question type and answer position on the test set of ELITR-Bench-QA in single-turn mode. The reported numbers correspond to the average scores from 1 to 10 (higher is better) obtained by a GPT-4 evaluator. The number N below a subset indicates the corresponding subset size. Boldface numbers correspond to the best performance among proprietary or open models.
Model Question type Answer position
@@
GPT-3.5 7.91 6.94 7.68 7.79 7.33 7.45 7.76 7.37
GPT-4 8.56 8.29 8.28 8.29 8.36 8.29 8.32 8.57
GPT-4o 8.68 8.12 8.60 8.92 8.17 8.67 8.42 8.56
LongAlpaca-7B 5.35 5.37 6.35 6.79 5.81 5.80 4.97 5.53
LongAlpaca-13B 7.19 5.47 6.47 6.00 5.93 5.95 6.85 6.59
LongChat-7B-v1.5 6.88 4.94 6.33 4.17 6.41 4.91 5.89 5.77
Vicuna-7B-v1.5 6.13 5.65 5.40 2.88 5.89 5.21 4.96 6.12
Vicuna-13B-v1.5 6.96 6.68 5.48 5.54 6.35 6.41 6.55 6.87
LongAlign-7B 6.93 6.33 6.00 5.88 7.09 6.39 6.47 5.66
LongAlign-13B 6.08 6.74 5.97 5.75 6.71 6.21 6.33 5.95
LLaMA-3.1-8B 8.18 7.53 7.53 8.67 7.95 7.60 8.00 7.77
Phi-3-small 7.67 6.78 7.85 8.25 7.57 7.36 7.06 7.22
In this section, we provide the full results split by question type and answer position obtained on ELITR-Bench-QA’s test set in the single-turn setting. The results are reported in Table 3. Looking at the global model performance over the different question types and answer positions, we do not identify any clear trend highlighting a question type or position answer as notably easier or harder. However, the Who questions seem to be on average slightly easier to answer. In contrast, the What questions were comparatively more challenging than other types for the best performing models (GPT-3.5, GPT-4, GPT-4o, LLaMA-3.1-8B, and Phi-3-small). This is not surprising as What questions sometimes require complex answers that go beyond simply listing entities, dates or numbers. Interestingly, LLaMA-2-based models struggled the most with the How many questions. Although the amount of such questions is very limited (8 in the test set) which calls for caution on tentative interpretations, this seems to suggest that LLaMA-2 models are notably less proficient at dealing with quantities and numbers than GPT models and more recent open LLMs such as LLaMA-3.1 and Phi-3.
Regarding the answer position, we did not notice any strong “lost in the middle” effect [31] which posits that information located in the middle section of long contexts is harder to access for LLMs. We verified this by conducting a statistical significance test on the scores obtained by each individual model, whose details are described in Appendix C.3.
5.3 Robustness to transcript noise
In Section 5.1, we studied how long-context LLMs fare at answering questions when having access to relatively clean meeting transcripts. However, in many practical scenarios, the quality of the transcripts might be degraded due to different factors: e.g., the audio recording conditions, the presence of accented speech, or simply the lacking capabilities of the ASR model. We then sought to understand how robust long-context LLMs are in the presence of a noisy transcript. For that purpose, we tested the three best-performing models from Table 2 – i.e, GPT-4o, [^12] Given the similar performance of GPT-4 and GPT-4o, we only retained GPT-4o due to its lower cost. LLaMA-3.1-8B, and Phi-3-small – on the test set of ELITR-Bench-QA in single-turn mode, using transcripts with varied levels of noise. The noisy transcripts were obtained following the procedure introduced in Section 3 and detailed in Appendix A. The results are reported in Fig. 1. In this experiment, a single seed is used to limit the cost incurred by GPT-4-based evaluation.
We observe in Fig. 1 that while the gap between the two open models and GPT-4o is small on the clean transcript (around 1 point), it widens significantly as the noise level is increased. Interestingly, GPT-4o also seems to resists much better to mild noise (0.2 and 0.4) in comparison to other models. Even at very high noise levels (0.8 and 1.0), its average score remains above 6 which is similar to the performance of LLaMA-2-based models from Table 2. All in all, we conclude that while the most recent open long-context LLMs approach GPT-4/4o capabilities on clean transcripts, there remains an important gap when noisier context is used.
Figure 1: Comparison of the scores obtained for GPT-4o, LLaMA-3.1-8B, and Phi-3-small using transcripts with varied levels of noise on the test set of ELITR-Bench-QA in single-turn mode. Indicated levels of noise correspond to the target Word Error Rates set in our noise injection procedure.